# Generative Pricing of Basket Options via Signature-Conditioned Mixture Density Networks

### v1

@@ Abstract
We present a generative framework for pricing European-style basket options by learning the conditional terminal distribution of the log arithmetic-weighted basket return. A Mixture Density Network (MDN) maps time-varying market inputs—encoded via truncated path signatures—to the full terminal density in a single forward pass. Traditional approaches either impose restrictive assumptions or require costly re-simulation whenever inputs change, limiting real-time use. Trained on Monte Carlo (MC) under GBM with time-varying volatility or local volatility, the MDN acts as a reusable surrogate distribution: once trained, it prices new scenarios by integrating the learned density. Across maturities, correlations, and basket weights, the learned densities closely match MC (low KL) and produce small pricing errors, while enabling train-once, price-anywhere reuse at inference-time latency.
@@ 1.2 Mixture Density Networks for Distribution Approximation
MDNs have proven effective in modelling stochastic systems across diverse fields, from acoustic modelling [Zen_2014_MDNAcousticModeling] and bioinformatics [Ji_2005_BetaMixtureBioinfo] to probabilistic forecasting [Zhang_2020_RegionalWindMDN] and volatility estimation. Several studies have also explored the financial applications of MDNs. Schittenkopf and Dorffner [Schittenkopf_2001_RND_MDNet] extended the MDN framework to extract risk-neutral densities from observed option prices, while Schittenkopf et al. [Schittenkopf_1998_MDNVolatility] utilized MDNs to estimate and forecast stock market volatility. Li et al. [Li_2024_XRMDN] introduced a recurrent MDN architecture for demand forecasting in high-volatility mobility-on-demand systems.
@@ 1.3 Signature-Conditioned MDNs for Basket Option Pricing
Relation to Generative Models in Finance: The Mixture density network approach for conditional density approximation belongs to the broader class of generative models. MDNs directly model $p(y|\textbf{x})$ as a finite mixture and train via exact negative log-likelihood (NLL). In one-dimensional settings, this yields a lightweight, numerically stable, and inherently multi-modal representation. By contrast, variational autoencoders (VAEs) optimize an evidence-lower-bound (ELBO) surrogate [Kingma2014VAE, Rezende2014StochasticBackprop] and introduce additional encoder-decoder structures; while flexible, they add unnecessary latent complexity for 1D conditionals and are prone to posterior collapse. Normalizing flows provide exact likelihoods [Papamakarios2017MAF, Durkan2019NSF, Papamakarios2021NFJMLR] and excel in high-dimensional outputs, but are often over-parameterized for 1D tasks unless carefully constrained. Finally, diffusion and score-based models [Ho2020DDPM, Song2021SDE] offer state-of-the-art generative fidelity but require multi-step sampling and heavy conditioning, leading to higher computational cost. For the univariate conditional densities considered in this study, the MDN strikes the best balance between fidelity, numerical stability, and deployment cost. Nevertheless, flow- or diffusion-based frameworks remain promising candidates for future multivariate extensions, where richer joint structures among asset components become critical.
@@ 1.4 Background on Path Signatures
The signature of a path, introduced by Chen [chen_1957_IntgofPaths, chen_1977_ItePathIntg], is defined as an infinite sequence of iterated integrals of the path over increasing tensor orders. In practice, one often works with a truncated signature, and this truncation is theoretically justified by the fact that any path of finite $p$-variation is uniquely determined by its signature truncated at order $\lfloor p\rfloor$ (see Chapter 7 of [Friz_2010_RoughPaths]). Its ability to capture the essential characteristics of time series in a non-parametric manner has led to successful applications in various domains. In financial time series analysis, signatures have been effectively employed for both classification and prediction tasks [GyurkoLyons_2014_Extracting_signature_financial_data, LyonsNi_2014_feature_set_financialdata], demonstrating their utility in handling complex, high-frequency data. Beyond predictive modelling, signatures have been applied to solve forward-backward stochastic differential equations (FBSDEs) numerically [QiFeng_2023_DeepSignatureFBSDEAlgo], particularly in non-Markovian settings where path-dependence plays a critical role.
@@ 2.2 Mixture Density Network (Maximum Likelihood)
  • Mixing coefficients $\pi(\mathbf{x})=(\pi_{1}(\mathbf{x}),\cdots,\pi_{d}(\mathbf{x})),$
  • Means $\mu(\mathbf{x})=(\mu_{1}(\mathbf{x}),\cdots,\mu_{d}(\mathbf{x})),$
  • Standard deviations $\delta(\mathbf{x})=(\delta_{1}(\mathbf{x}),\cdots,\delta_{d}(\mathbf{x})),$
@@
  • Standard deviations $\delta_{j}(\mathbf{x})$ must be strictly positive. A common choice is the exponential function [Bishop_1994_MDN] to map unconstrained outputs to positive values, though alternatives such as ELU [Normandin_2023_LinearPretrainRMDN, Wu_2022_3DHumanPoseMDN, Yang_2021_GANMDNInverseModeling] and softplus [Goodfellow_2016_DeepLearning, Han_2022_SurvivalMDN, Hepp_2022_MDNIntervals, Muzakka_2024_RBSMDN, Razavi_2024_FRMDN, Wang_2022_LikelihoodFreeMDN] are sometimes preferred to improve numerical stability.
@@ 4.1 Basket of Assets Following GBM with Time-varying Volatility
We consider a basket of two assets ($N=2$), where each asset has equal weight $(w_{1}=w_{2}=0.5)$. For the simulation of the training dataset, we set the initial asset price as $S(0)=(S_{1}(0),S_{2}(0))=(1.0,1.0)$. The time-varying parameters $r(t)$, $q(t)$ and $\sigma(t)$ are simulated using the CIR model, with the corresponding parameter values and initial ranges summarized in Table 1.
@@ 4.1.4 Numerical Results
The time series of the risk-free interest rate, dividend yields, and volatilities, along with the correlation matrix, serve as inputs to the MDN. Assume, Figure 3a presents the time series of the risk-free interest rate, along with the dividend yields and volatilities for the two assets in the basket, simulated over a one-year horizon (252 trading days). The correlation structure among these assets is specified by the correlation matrix
@@ 4.2.5 Numerical Results
Figure 7a shows the time series of the risk-free interest rate and the dividend yields for the two assets in the basket. Using this market condition, we analyze three different combinations of the basket weights. The resulting distributions of the basket’s weighted returns, estimated by the mixture density network (MDN), are compared with those obtained via Monte Carlo (MC) simulation in Figures 7b, 7c and 7d, each corresponding to different weight vectors. These comparisons are carried out across multiple maturities. For each case, the Kullback–Leibler (KL) divergence between the MDN and MC distributions is reported alongside the corresponding maturity, denoted by ”D_KL”.
@@
The risk-free interest rate and dividend yields for two assets are now given in Figure 9a. For three different sets of basket weights, Figures 9b, 9c, and 9d present a comparison between the MDN-estimated and Monte Carlo-simulated distributions of the weighted basket return across various maturities. The associated Kullback–Leibler (KL) divergence between the two distributions is also reported for each case.
The resulting relative percentage errors in European option prices, corresponding to the basket weights used in Figure 9b and 9c, are reported in Figure 10 and  11 respectively.
These numerical results demonstrate that the MDN can successfully learn the conditional distribution based on the parameters of the local volatility function. Notably, the MDN is also parametrized by the basket weights, providing additional flexibility: if the basket composition changes, the trained model remains applicable for pricing options on the new weighted return, eliminating the need to retrain a separate model.

### v2

# arXiv:2511.09061v2[q-fin.PR] 27 Feb 2026

@@ Abstract
We present a generative framework for pricing European-style basket options by learning the conditional terminal distribution of the log arithmetic-weighted basket return. A Mixture Density Network (MDN) maps time-varying market inputs encoded via truncated path signatures to the full terminal density in a single forward pass. Traditional approaches either impose restrictive assumptions or require costly re-simulation whenever inputs change, limiting real-time use. Trained on Monte Carlo (MC) under GBM with time-varying volatility or local volatility, the MDN acts as a reusable surrogate distribution: once trained, it prices new scenarios by integrating the learned density. Across maturities, correlations, and basket weights, the learned densities closely match MC (low KL) and produce small pricing errors, while enabling train-once, price-anywhere reuse at inference-time latency.
@@ 1.2 Mixture Density Networks for Distribution Approximation
MDNs have proven effective in modelling stochastic systems across diverse fields, from acoustic modelling [28] and bioinformatics [12] to probabilistic forecasting [29] and volatility estimation. Several studies have also explored the financial applications of MDNs. Schittenkopf and Dorffner [23] extended the MDN framework to extract risk-neutral densities from observed option prices, while Schittenkopf et al. [22] utilized MDNs to estimate and forecast stock market volatility. Li et al. [14] introduced a recurrent MDN architecture for demand forecasting in high-volatility mobility-on-demand systems.
@@ 1.3 Signature-Conditioned MDNs for Basket Option Pricing
Relation to Generative Models in Finance: The Mixture density network approach for conditional density approximation belongs to the broader class of generative models. MDNs directly model $p(y|\textbf{x})$ as a finite mixture and train via exact negative log-likelihood (NLL). In one-dimensional settings, this yields a lightweight, numerically stable, and inherently multi-modal representation. By contrast, variational autoencoders (VAEs) optimize an evidence-lower-bound (ELBO) surrogate [13, 21] and introduce additional encoder-decoder structures; while flexible, they add unnecessary latent complexity for 1D conditionals and are prone to posterior collapse. Normalizing flows provide exact likelihoods [19, 4, 18] and excel in high-dimensional outputs, but are often over-parameterized for 1D tasks unless carefully constrained. Finally, diffusion and score-based models [11, 24] offer state-of-the-art generative fidelity but require multi-step sampling and heavy conditioning, leading to higher computational cost. For the univariate conditional densities considered in this study, the MDN strikes the best balance between fidelity, numerical stability, and deployment cost. Nevertheless, flow- or diffusion-based frameworks remain promising candidates for future multivariate extensions, where richer joint structures among asset components become critical.
@@ 1.4 Background on Path Signatures
The signature of a path, introduced by Chen [2, 3], is defined as an infinite sequence of iterated integrals of the path over increasing tensor orders. In practice, one often works with a truncated signature, and this truncation is theoretically justified by the fact that any path of finite $p$-variation is uniquely determined by its signature truncated at order $\lfloor p\rfloor$ (see Chapter 7 of [6]). Its ability to capture the essential characteristics of time series in a non-parametric manner has led to successful applications in various domains. In financial time series analysis, signatures have been effectively employed for both classification and prediction tasks [8, 15], demonstrating their utility in handling complex, high-frequency data. Beyond predictive modelling, signatures have been applied to solve forward-backward stochastic differential equations (FBSDEs) numerically [5], particularly in non-Markovian settings where path-dependence plays a critical role.
@@ 2.2 Mixture Density Network (Maximum Likelihood)
  • Mixing coefficients, $\pi(\mathbf{x})=(\pi_{1}(\mathbf{x}),\cdots,\pi_{d}(\mathbf{x})),$
  • Means, $\mu(\mathbf{x})=(\mu_{1}(\mathbf{x}),\cdots,\mu_{d}(\mathbf{x})),$
  • Standard deviations, $\delta(\mathbf{x})=(\delta_{1}(\mathbf{x}),\cdots,\delta_{d}(\mathbf{x})),$
@@
  • Standard deviations $\delta_{j}(\mathbf{x})$ must be strictly positive. A common choice is the exponential function [1] to map unconstrained outputs to positive values, though alternatives such as ELU [17, 26, 27] and softplus [7, 9, 10, 16, 20, 25] are sometimes preferred to improve numerical stability.
@@ 4.1 Basket of Assets Following GBM with Time-varying Volatility
We consider a basket of assets, where each asset has a predefined weight in the basket. The time-varying parameters $r(t)$, $q(t)$ and $\sigma(t)$ are simulated using the CIR model, with the corresponding parameter values and initial ranges summarized in Table 1.
@@ 4.1.2 MDN Architecture:
where $\varepsilon^{0}$ is a small positive constant used to ensure numerical stability in the standard deviation output. Unless
@@ 4.1.4 Numerical Results
2-Asset Basket: We consider a basket of two assets ($N=2$), where each asset has equal weight $(w_{1}=0.5,w_{2}=0.5)$. For the simulation of the training dataset, we set the initial asset price as $S(0)=(S_{1}(0),S_{2}(0))=(1.0,1.0)$. The time series of the risk-free interest rate, dividend yields, and volatilities, along with the correlation matrix, serve as inputs to the MDN. Assume, Figure 3a presents the time series of the risk-free interest rate, along with the dividend yields and volatilities for the two assets in the basket, simulated over a one-year horizon (252 trading days). The correlation structure among these assets is specified by the correlation matrix
@@
3-Asset Basket: We consider a basket of three assets ($N=3$), where assets has weight $(w_{1}=0.33,w_{2}=0.33,w_{3}=0.34)$. For the simulation of the training dataset, we set the initial asset price as $S(0)=(S_{1}(0),S_{2}(0),S_{3}(0))=(1.0,1.0,1.0)$. We have used almost the same MDN structure as mentioned in Section 4.1.2, but with $g_{\mu}(\mathbf{x})=\mathbf{x}+\tanh{\mathbf{x}}$. Figure 7a presents the time series of the risk-free interest rate, along with the dividend yields and volatilities for the three assets in the basket. The correlation structure among these assets is specified by the correlation matrix
$\mathbf{R}=\begin{pmatrix}1&-0.9968&0.9770\\ -0.9968&1&-0.9862\\ 0.9770&-0.9862&1\end{pmatrix}$.
Then, Figure 7b compares the corresponding MDN and MC return distributions, with the associated KL divergences. The resulting relative percentage pricing errors for European options are reported in Figure 8. These results are shown separately for European call and put options at various strikes and maturities, offering insights into the pricing accuracy achieved by the MDN approximation.
In a second scenario, the correlation matrix is changed to
$\mathbf{R}=\begin{pmatrix}1&0.5138&-0.2963\\ 0.5138&1&0.6030\\ -0.2963&0.6030&1\end{pmatrix}$,
4-Asset Basket: We consider a basket of four assets ($N=4$), where each assets has same weight $(w_{1}=0.25,w_{2}=0.25,w_{3}=0.25,w_{4}=0.25)$. For the simulation of the training dataset, we set the initial asset price as $S(0)=(S_{1}(0),S_{2}(0),S_{3}(0),S_{4}(0))=(1.0,1.0,1.0,1.0)$. We have used the same MDN structure as the 3-Asset Basket case. Figure 11a presents the time series of the risk-free interest rate, along with the dividend yields and volatilities for the three assets in the basket. The correlation structure among these assets is specified by the correlation matrix
$\mathbf{R}=\begin{pmatrix}1&-0.6663&1&0.1460\\ -0.6663&1&-0.6652&0.0308\\ 1&-0.6652&1&0.1386\\ 0.1460&0.0308&0.1386&1\end{pmatrix}$.
Then, Figure 11b compares the corresponding MDN and MC return distributions, with the associated KL divergences. The resulting relative percentage pricing errors for European options are reported in Figure 12. These results are shown separately for European call and put options at various strikes and maturities, offering insights into the pricing accuracy achieved by the MDN approximation.
In a second scenario, the correlation matrix is changed to
$\mathbf{R}=\begin{pmatrix}1&-0.9780&-0.8317&0.8888\\ -0.9780&1&0.9164&-0.9471\\ -0.8317&0.9164&1&-0.8868\\ 0.8888&-0.9471&-0.8868&1\end{pmatrix}$.
@@ 4.2.5 Numerical Results
Figure 15a shows the time series of the risk-free interest rate and the dividend yields for the two assets in the basket. Using this market condition, we analyze three different combinations of the basket weights. The resulting distributions of the basket’s weighted returns, estimated by the mixture density network (MDN), are compared with those obtained via Monte Carlo (MC) simulation in Figures 15b, 15c and 15d, each corresponding to different weight vectors. These comparisons are carried out across multiple maturities. For each case, the Kullback–Leibler (KL) divergence between the MDN and MC distributions is reported alongside the corresponding maturity, denoted by ”D_KL”.
@@
The risk-free interest rate and dividend yields for two assets are now given in Figure 17a. For three different sets of basket weights, Figures 17b, 17c, and 17d present a comparison between the MDN-estimated and Monte Carlo-simulated distributions of the weighted basket return across various maturities. The associated Kullback–Leibler (KL) divergence between the two distributions is also reported for each case.
The resulting relative percentage errors in European option prices, corresponding to the basket weights used in Figure 17b and 17c, are reported in Figure 18 and  19 respectively. These numerical results demonstrate that the MDN can successfully learn the conditional distribution based on the parameters of the local volatility function. Notably, the MDN is also parametrized by the basket weights, providing additional flexibility: if the basket composition changes, the trained model remains applicable for pricing options on the new weighted return, eliminating the need to retrain a separate model.